Determine ‘a’ and ‘b’ as to make the cylinder y2= 4ax orthogonal to the ellipsoid at the point (1, 2, 1).
Find the directional derivative of f(x,y)=2x^2+y^2 at the point (-2,3) in the direction of u ⃗=4i ̂-5j ̂.

Show that the function f(x,y)=e^x siny+e^y cosx satisfies the Laplace’s equation
1
Expert's answer
2021-05-07T10:31:51-0400

Ans:-

"f(x,y)=e^x siny+e^y cosx"

partial derivative with respect to x partial derivative with respect to y

"\frac{\partial f}{\partial x} =e^x siny-e^ysinx" , "\frac{\partial f}{\partial y}=e^xcosy+e^ycosx"


again partial derivative with respect to x again partial derivative with respect to y

"\frac{\partial^2 f}{\partial x^2}=e^xsiny-e^ycosx" "-(i)" "\frac{\partial^2 f}{\partial x^2}=-e^xsiny+e^ycosx" "-(ii)"


Add these two equations

"\Rightarrow \frac{\partial^2 f}{\partial x^2} +\frac{\partial^2 f}{\partial y^2}=e^xsiny-e^ycosx+(-e^xsiny+e^ycosx)" "=0"


"\Rightarrow \frac{\partial^2 f}{\partial x^2} +\frac{\partial^2 f}{\partial y^2}=0"

Hence Laplace's equation will be satisfied.